What is the sum of the positive prime factors of 70?
Found 2 solutions by stanbon, jim_thompson5910:
Answer by stanbon(75887)

What is the sum of the positive prime factors of 70?
---------------------
The prime factors are 7,2,5
The sum is 7+2+5 = 14
=======================
Cheers,
Stan H.

Answer by jim_thompson5910(35256)

The number 70 has the prime factorization: 70 = 2*5*7


Simply add these factors to get 2+5+7=14